OpenStudy (anonymous):

A corner store sells two kinds of baked goods: cakes and pies. A cake costs $5 and a pie costs $7. In one day, the store sold 15 baked goods for a total of $91. How many cakes did they sell?

OpenStudy (callisto):

let x and y be the no of cakes and no. of pies sold respectively x+y=15 -> y=15-x (1) 5x+7y = 91 - (2) put (1) into (2) 5x + 7(15-x) = 91 105 -2x = 91 2x = 17 x=7 therefore 7 pieces of cake are sold